# AST-T5: Structure-Aware Pretraining for Code Generation and Understanding

### v1

@@ A.4 Pretraining Hyperparameters
A.5 Evaluation Benchmarks
Table 4: Overview of our evaluation benchmarks detailing test set size, task type, and evaluation metric for each task. For MBPP, we follow Nijkamp et al. (2023) and evaluate our model on the entire “sanitized” subset without few-shot prompts. For evaluation metrics, “Pass@1” indicates code execution on unit-tests provided in the benchmark using a single generated code per example, with reported pass rates. “Exact Match” evaluates textual equivalence without execution by comparing two canonicalized code pieces. We omit “BLEU scores” because high BLEU values (>> 50) can still correspond to unexecutable or significantly flawed code (Lu et al., 2021), which is not useful in real-world applications. We also discuss evaluation results using the CodeBLEU (Ren et al., 2020) metric in Section A.7.
Size Type Metric
HumanEval 164 Text-to-Code Generation Pass@1
MBPP 427 Text-to-Code Generation Pass@1
Concode 2,000 Text-to-Code Generation Exact Match
Bugs2Fix 12,379 Code-to-Code Transpilation Exact Match
Java-C# 1,000 Code-to-Code Transpilation Exact Match
BigCloneBench 415,416 Code Understanding F1
Defect Detection 27,318 Code Understanding Accuracy
A.6 Evaluation Results on Multi-Lingual Code Generation
@@
Table 5 presents a comparative analysis of our AST-T5 model on Python and Java subsets of the multi-lingual HumanEval and MBXP benchmarks (Athiwaratkun et al., 2023). This analysis includes models such as BLOOM (BigScience, 2021), OPT (Zhang et al., 2022), and various configurations of CodeGen (Nijkamp et al., 2023), as reported in Athiwaratkun et al. (2023). Our results show AST-T5’s superior performance across all benchmarks compared to the CodeGen-multi-350M. Notably, although CodeGen-mono-350M, tailored for Python, surpasses AST-T5 in the MBPP benchmark, it significantly underperforms in the Java subset. Furthermore, AST-T5, having 226M parameters, outperforms larger counterparts like BLOOM-7.1B and OPT-13B.
A.7 Evaluation Results in CodeBLEU

### v2

@@ A.3 Pretraining Hyperparameters
A.4 Evaluation Results on Multi-Lingual Code Generation
@@
Table 5 presents a comparative analysis of our AST-T5 model on Python and Java subsets of the multi-lingual HumanEval and MBXP benchmarks (Athiwaratkun et al., 2023). This analysis includes models such as BLOOM (BigScience, 2021), OPT (Zhang et al., 2022), and various configurations of CodeGen (Nijkamp et al., 2023), as reported in Athiwaratkun et al. (2023). Our results show AST-T5’s superior performance across all benchmarks compared to the CodeGen-multi-350M. Furthermore, AST-T5, having 226M parameters, outperforms larger counterparts like BLOOM-7.1B and OPT-13B.
A.5 Evaluation Results in CodeBLEU